A study was conducted on a high-pressure inlet fogging method for a gas turbine engine. Researchers considered two models. Partial output is provided below.

Source

Df

SS

MS

F

Regression

5

148526859

29705372

94

Error

61

19370350

317547

Total

???

167897208

Source

Df

SS

MS

F

Regression

???

142586570

47528857

???

Error

63

25310639

401756

Total

???

167897208

Answer the following questions rounding off at the nearest integer.

The number of observations collected for the study is equal to

The number of predictors considered in the reduced model is equal to

The F-test for global utility of the reduced model is equal to

Results suggest that the complete model provides additional information compared to the reduced model. In fact, the F-test for nested models is equal to  which is greater than the rejection region F(alpha =0.05) with numerator degrees of freedom equal to  and denominator degrees of freedom equal to  .
Answer #1

The number of observations collected for the study is equal to =5+61+1 =67

The number of predictors considered in the reduced model is equal to =142586570/47528857 =3

The F-test for global utility of the reduced model is equal to =47528857/401756 =118.30 ~118

Partial F test statistic =((SSEr-SSEc)/(c-r))/(SSEc/(n-c-1)) = 9.353409
numerator df =(c-r) = 2
and denominator df =(n-c-1) = 61

  In fact, the F-test for nested models is equal to 9 which is greater than the rejection region F(alpha =0.05) with numerator degrees of freedom 2 equal to  and denominator degrees of freedom equal to 61 .